# supSupplementary References

### v1

@@ Abstract
Systems that answer questions by reviewing the scientific literature are becoming increasingly feasible. To draw reliable conclusions, these systems should take into account the quality of available evidence, placing more weight on studies that use a valid methodology. We present a benchmark for measuring the methodological strength of biomedical papers, drawing on the risk-of-bias framework used for systematic reviews. The four benchmark tasks, drawn from more than 500 papers, cover the analysis of research study methodology, followed by evaluation of risk of bias in these studies. The benchmark contains 2000 expert-generated bias annotations, and a human-validated pipeline for fine-grained alignment with research paper content. We evaluate a range of large language models on the benchmark, and find that these models fall significantly short of expert-level performance. By providing a standardized tool for measuring judgments of study quality, the benchmark can help to guide systems that perform large-scale aggregation of scientific data. The dataset is available at https://github.com/RoBBR-Benchmark/RoBBR.
@@ 1 Introduction
Systems that automatically answer questions by reviewing the scientific literature are becoming increasingly feasible. These systems offer the potential to provide scientists with on-demand access to knowledge that is synthesized from across the literature. This can help scientists understand what is already known about topics outside of their research focus, and help clinicians and practitioners keep up to date with best practices.
When assessing what is known about a field, not all studies should be weighed equally I et al. [2023]. Studies with stronger methodologies contribute more to a body of evidence than those with weaker methodologies. By weighing studies appropriately, systems can increase the reliability of their summaries and recommendations Turner et al. [2009].
In biomedical research, there is a large body of work which investigates the factors that decrease the validity of a study’s methods, and best practices for addressing these issues I et al. [2023], Welton et al. [2009]. Reporting bias occurs when there are systematic differences in how outcomes are reported or disclosed between the groups that are compared (e.g., the results of the treatment group are more frequently or favorably published than those in the placebo group). This bias is often best addressed by adhering to strict reporting standards and protocols, such as registering the study in advance and committing to publish all results Kotz and West [2022]. Other biases, such as detection bias, attrition bias, and selection biases, can frequently be addressed using careful study design and implementation Berger [2005], JPT et al. [2023], I et al. [2023].
This paper presents the RoBBR benchmark for assessing the methodological strength of biomedical studies. The benchmark follows the risk of bias framework developed by Cochrane Jpt et al. [2023], JPT et al. [2023], Higgins et al. [2011], Practice and of Care [EPOC], Sterne et al. [2014], a global independent network that conducts systematic reviews of healthcare interventions. Cochrane’s risk of bias framework is widely recognized as a standard for evaluating the quality and reliability of research studies. It serves as an important resource for national health agencies in various countries, including the United States Viswanathan et al. [2012], the United Kingdom Alderson and Tan [2011], and Australia NHMRC [2019]. Notably, Cochrane’s systematic reviews are used to formulate national and international guidelines on healthcare practices Bunn et al. [2015]. Furthermore, empirical comparisons between Cochrane and non-Cochrane systematic reviews consistently demonstrate higher methodological quality in those conducted by Cochrane, highlighting its critical role in advancing evidence-based healthcare Goldkuhle et al. [2018], Page et al. [2018], Matthias et al. [2020], Tsoi et al. [2020].
Figure 1: The figure illustrates the components of a risk of bias assessment. Given a bias category and a research study, a meta-reviewer identifies factors that make the study high or low risk for that bias. The meta-reviewer summarizes their findings in a support judgment, and then make a final decision about the risk level.
@@ 2 Background
The RoBBR benchmark aims to provide an expert-validated tool for assessing whether models can make risk of bias judgments for biomedical papers. The benchmark uses a risk-of-bias framework developed in the context of systematic reviews, which is the gold standard for assessing biomedical studies. We provide more background on systematic reviews and the risk of bias framework below.
@@
Systematic reviews and meta-analyses are tools used in evidence-based medicine, which synthesize evidence from multiple studies to answer important clinical questions. They help researchers and healthcare professionals make informed decisions based on the best available evidence Ahn and Kang [2018], TJ et al. [2023], JJ et al. [2023].
A systematic review is a comprehensive search of the literature to identify all relevant studies on a particular topic TJ et al. [2023], C et al. [2023]. The review process follows a pre-specified protocol to minimize bias and ensure transparency. The protocol outlines the research question, search strategy, study selection criteria, data extraction methods, and plans for quality assessment and synthesis of result Jpt et al. [2023], Page et al. [2021].
A meta-analysis is a type of systematic review that uses statistical methods to combine the results from multiple studies Ahn and Kang [2018]. By pooling the data across studies, a meta-analysis can provide a more precise estimate of the treatment effect JJ et al. [2023], Guevara et al. [2004] and explore potential sources of heterogeneity Higgins and Thompson [2002].
The research question for a systematic review is often structured using the PICO framework, which stands for Participant, Intervention, Comparison, and Outcome Aslam and Emmanuel [2010]. The PICO elements help define the key components of the research question and guide the study selection process J et al. [2023], JE et al. [2023a]. For example, a systematic review on the effectiveness of a new hypertension drug might have the following PICO: Participant: Adults with primary hypertension; Intervention: New antihypertensive drug; Comparison: Placebo or standard care; Outcome: Change in systolic and diastolic blood pressure.
By clearly defining the PICO elements, the review authors can develop a focused search strategy and establish explicit criteria for selecting studies to include in the review J et al. [2023], JE et al. [2023b].
Inclusion criteria typically specify the characteristics of the study population, the interventions and comparisons of interest, the outcomes to be measured, and the study designs to be included JE et al. [2023b], Page et al. [2021], Tawfik et al. [2019] (e.g., randomized controlled trials). Exclusion criteria describe reasons why a study must be excluded from an analysis, such as using the wrong study design C et al. [2023], Page et al. [2021], Tawfik et al. [2019].
@@
Risk of bias assessment involves evaluating each included study to determine the likelihood that its results may be biased JPT et al. [2023], Practice and of Care [EPOC], Sterne et al. [2014]. The Cochrane tool for assessing risk of bias Higgins et al. [2011] is widely used and assesses several types of bias, including random sequence generation, allocation concealment, blinding of participants and outcome assessors, incomplete outcome data, selective reporting, and other sources of bias. During the assessment, two review authors independently evaluate the risk of bias using the Cochrane tool and resolve disagreements through discussion. They record their rationale for each bias assessment. These rationales are referred to as "support judgments."
For example, a support judgment of Reporting Bias could be "The study protocol was registered (ClinicalTrials.gov NCT01863706) but not all of the outcomes projected by methodological descriptions were reported as results in the study report (cases of diarrhoea, nausea and vomiting were not completely reported). Moreover, the study publication reports outcomes (hypotension, nausea, transfusion) not listed in the registered protocol." A decision regarding the risk of reporting bias is then made on the basis of the support judgment. See Figure 1 for an illustration.
Studies with a high risk of bias may overestimate or underestimate the true treatment effect, leading to inaccurate conclusions in the systematic review. In this paper, we consider eight categories of biases: selection bias, attrition bias, performance bias, detection bias, reporting bias, analysis bias, special bias (specific to individual meta-analyses), and other bias (i.e. biases not covered by the other categories). We benchmark model performance on assessing papers for each of these bias categories.
@@ 3 Benchmark Development
The RoBBR benchmark is derived from 63 Cochrane meta-analyses and 532 papers referenced in those meta-analyses. All of the meta-analyses and papers have a Creative Commons license or are public domain. The benchmark is designed to evaluate a model’s ability to judge the risk of bias in biomedical reports.
The data were collected in April 2024. We included all meta-analyses from Cochrane that are in the BioC database, are classified as intervention reviews, have a CC-BY-NC license, and contribute at least one data point for each task. We obtained the XML format of the papers included in these meta-analyses either through the BioC API or by manually downloading and converting PDFs to XML using GROBID. We then extract task data from the XML of these papers.
The benchmark is divided into four main tasks (See in Appendix A.3 for detailed statistics):
  • Study Inclusion/Exclusion (Task 1): Determine whether a study’s objectives and methodology fit a set of requirements.
  • Bias Retrieval (Task 2): Retrieve sentences from the paper that support a judgment about the study’s risk of bias.
  • Support Judgment Selection (Task 3): From a list of options, select the judgment that best supports a determination of the study’s risk of bias.
  • Risk Level Determination (Task 4): Assess the study’s risk of specific biases.
These four tasks mirror the procedure followed by meta-reviewers. Initially, reviewers decide if a study should be included or excluded based on predefined criteria, which reflect the goals of the review (Task 1). Included studies will be analyzed for indications of the risk of bias (Task 2). Reviewers will then summarize the evidence collected from the study, and analyze its implications (Task 3). Finally, they assess the risk level of specific biases for each identified bias (Task 4).
By structuring the dataset around these tasks, we provide a comprehensive framework for simulating the nuanced process of risk of bias assessment in meta-analysis.
3.1 Train/Test Split
We randomly select 23 meta-analyses and their 228 corresponding papers in the dev set. We put the other 40 meta-analyses and their corresponding 304 papers in the test set.
Figure 2: Study Inclusion/Exclusion. In Task 1, the model sees a search protocol and objective, and is asked to determine whether a study’s methodology fits those criteria.
3.2 Task 1: Study Inclusion/Exclusion
For each potentially relevant study to be included in a meta-analysis, the search protocol information and meta-analysis objective are provided as input, and the task is to decide if the study should be included in the meta-analysis. The objective is included in order to provide similar context to what is normally available for a meta-reviewer and make it easier to interpret the criteria in the search protocol. An example is provided in Figure 2.
3.3 Task 2: Risk of Bias Sentence Retrieval
When judging the risk of bias for a specific study, the review authors provide support judgments which justify their risk of bias rating. These support judgments are grounded in specific aspects of the study, which the support judgment describes. Task 2 aims to test a model’s ability to retrieve the correct source for the support judgment from a paper’s text. The support judgments are intentionally omitted in the Bias Retrieval task. This exclusion simulates the systematic review process by challenging the model to independently generate these justifications. Including them would come close to providing the answer, undermining the purpose of the task. See Figure 3 for an illustration.
Some review authors directly quote sentences from the paper in their support judgment, allowing for straightforward extraction and matching of these sentences with the paper’s text. However, most reviewers prefer to paraphrase the sentences or engage in deeper analysis, synthesizing information from multiple parts of the paper.
To match these judgments with sentences from the paper, we use a three part annotation pipeline.
First, a support judgment often contains multiple pieces of information. We decompose the support judgment into distinct, non-overlapping pieces, each focusing on a specific aspect of the original support judgment.
Second, we exclude any part of the support judgment containing a negation or non-specific commentary, because these parts of the judgment cannot be matched to any sentence. The remaining parts of the support judgments generally contain specific claims about the study.
Third, these aspects are mapped to specific sentences within the paper, allowing us to pinpoint the exact sources of information that underlie each aspect of the judgment. A single aspect could potentially be supported by multiple sentences from the paper, and similarly a single sentence can support multiple aspects.
Each of these steps, including the aspect decomposition, uses GPT-4 (gpt-4-0125) OpenAI [2023]. The prompts can be found in the Appendix C.
We manually inspect the 2221 support judgment aspects generated in the second stage, to ensure that all aspects are related to specific information from the paper.
In the next section, we evaluate annotation quality by comparing to human judgments.
Figure 3: Risk of Bias: Sentence Retrieval. In Task 2, the goal is to retrieve sentences from the paper (the Candidate Pool) which support a risk of bias judgment. Performance is evaluated against a Cochrane review support judgment, which has been split into aspects.
3.3.1 Evaluating LLM Annotations
To match each aspect from a support judgment with specific sentences in the paper, we need to evaluate each sentence (on average, 200 sentences) against each aspect. For a support judgment with 5 aspects, this results in approximately 1,000 decisions. Given the scale of our entire dataset, which includes over 500 (bias, study) pairs, this translates to more than 1 million decisions – an impractical task for human annotators to perform manually.
We hypothesize that determining whether a sentence matches an aspect is straightforward enough for GPT-4 to achieve human-level performance with a specialized prompt. Following the protocol introduced by Wang et al. [2023], we created a development set of 30 papers. A single support judgment aspect was chosen for each paper, and every sentence in the paper was annotated for that aspect, totaling 8789 decisions. We optimized the instruction prompt for GPT-4 on this development set. Details about the prompt and experimental procedure are provided in the Appendix C.
Once the prompt was finalized, four annotators were tasked with annotating 50 randomly sampled (aspect, full paper) pairs (all annotators saw the same pairs). Every sentence in each paper was annotated, requiring 13575 annotations in total. The annotators were divided into two teams, with each team consisting of two annotators. Each annotator performed the annotation tasks individually given a pre-defined annotation guideline. Annotators from the same team then collaborated to resolve differences and eliminate mistakes. Refer to the Appendix B for the full annotation guidelines.
Each team produced a set of annotation results. GPT-4 annotated the same (aspect, paper) pairs. We calculated Binary Accuracy, F1-binary, Spearman Correlation, and Kappa Coefficient between the two human teams and between each human team and GPT-4. Table 1 shows that the human teams matched each other and GPT-4 more than 99% of the time. The other agreement metrics have values in the low-to-mid 70’s, indicating substantial agreement among annotators. [^1] These metrics are lower than exact agreement because of class imbalance. Positive matches between sentences and aspects are rare. Bootstrapped hypothesis tests do not demonstrate any statistical difference between GPT-4 and the human annotators.
Table 1: Inter-annotator agreement among the two teams of humans and the human teams and GPT-4.
Metrics Human & Human Average Human & GPT Average p-value

### v2

# Measuring Risk of Bias in Biomedical Reports: The RoBBR Benchmark

@@ Abstract
Systems that answer questions by reviewing the scientific literature are becoming increasingly feasible. To draw reliable conclusions, these systems should take into account the quality of available evidence from different studies, placing more weight on studies that use a valid methodology. We present a benchmark for measuring the methodological strength of biomedical papers, drawing on the risk-of-bias framework used for systematic reviews. Derived from over 500 biomedical studies, the three benchmark tasks encompass expert reviewers’ judgments of studies’ research methodologies, including the assessments of risk of bias within these studies. The benchmark contains a human-validated annotation pipeline for fine-grained alignment of reviewers’ judgments with research paper sentences. Our analyses show that large language models’ reasoning and retrieval capabilities impact their effectiveness with risk-of-bias assessment. The dataset is available at https://github.com/RoBBR-Benchmark/RoBBR.
Measuring Risk of Bias in Biomedical Reports: The RoBBR Benchmark
Jianyou Wangthanks: Equal contribution   Weili Cao11footnotemark: 1   Longtian Bao   Youze Zheng   Gil Pasternak Kaicheng WangXiaoyue WangRamamohan PaturiLeon Bergen Laboratory for Emerging Intelligence University of California, San Diego {jiw101, w2cao, rpaturi, lbergen}@ucsd.edu
Figure 1: Reviewer finds Sentence 4, 66, 73, 183. The reviewer writes a support judgment based on these sentences. The support judgment indicates this biomedical study has a high risk for selective reporting bias. Both support judgment and the high risk rating are included in the systematic review.
@@ 1 Introduction
Systems that automatically answer questions by reviewing the scientific literature are becoming increasingly feasible. These systems, such as "Deep Research" offerings from OpenAI, Google, and Anthropic, have the potential to provide scientists with on-demand access to knowledge that is synthesized from across the literature. This can help scientists understand what is already known about topics outside of their research focus, and help clinicians and practitioners keep up to date with best practices.
When assessing what is known about a field, not all studies should be weighed equally Boutron et al. (2023). Studies with stronger methodologies contribute more to a body of evidence than those with weaker methodologies. By weighing studies appropriately, systems can increase the reliability of their summaries and recommendations Turner et al. (2009).
In biomedical research, there is a large body of work which investigates the factors that decrease the validity of a study’s methods, and best practices for addressing these issues Boutron et al. (2023); Welton et al. (2009). For instance, reporting bias occurs when there are systematic differences in how outcomes are reported or disclosed between the groups that are compared (e.g., the results of the treatment group are more frequently or favorably published than those in the placebo group). This bias can be mitigated by registering the study in advance and committing to publish all results Kotz and West (2022).
We introduce the RoBBR benchmark for evaluating the methodological strength of biomedical studies, which is referred to as their risk-of-bias levels in the context of this paper. RoBBR’s main task involves labeling these risk-of-bias levels. Importantly, we also introduce two novel subtasks: support sentence retrieval (SSR) and support judgment selection (SJS). The SSR subtask is designed to test a model’s ability to identify support sentences — specific sentences within an entire paper that signal potential biases. We created the SSR dataset using our annotation pipeline. The SJS subtask evaluates a model’s capacity to synthesize this retrieved information through reasoning to form a well-supported judgment. Through our in-depth analysis, we find that the abilities to reason and, especially, to retrieve information are important milestones in improving models’ capacity to assess risk-of-bias levels.
@@ 2 Background
We provide more background on systematic reviews and the risk-of-bias guidelines.
@@
A systematic review is a method in evidence-based medicine that synthesizes evidence from multiple studies to answer important clinical questions. They help researchers and healthcare professionals make informed decisions based on the best available evidence Ahn and Kang (2018); Lasserson et al. (2023); Deeks et al. (2023).
Risk-of-bias assessment involves evaluating each study to determine the likelihood that its results may be biased Higgins et al. (2023); Practice and of Care (EPOC); Sterne et al. (2014). Studies with a high risk of bias may overestimate or underestimate the true treatment effect, leading to inaccurate conclusions in the systematic review.
RoBBR follows the risk-of-bias guideline developed by Cochrane Sterne et al. (2019); Higgins et al. (2023, 2011); Practice and of Care (EPOC); Sterne et al. (2014), a global independent network that conducts systematic reviews of healthcare interventions. Cochrane’s risk-of-bias guideline is widely recognized as a standard for evaluating the quality and reliability of research studies. It serves as an important resource for national health agencies in various countries, and are used to formulate national and international guidelines on healthcare practices Viswanathan et al. (2012); Alderson and Tan (2011); NHMRC (2019); Bunn et al. (2015).
@@
During reviewer’s assessment of a study’s risk of bias, two review authors independently evaluate the risk of bias using the Cochrane tool and resolve disagreements through discussion. They record their rationale for each bias assessment. These rationales are referred to as "support judgments".
Figure 2: Given a support judgment, our annotation pipeline produces the Aspect-to-Sentence Mapping, a many-to-many relationship. The mapping on the right-side shows Sentence 66 covers both Aspect 1 and 2, so if a model retrieves Sentence 66, it should not retrieve Sentence 4 to avoid redundancy.
3 Related Work
Risk-of-Bias Labeling
Previous research focuses on creating datasets and training models to label risk-of-bias levels Lai et al. (2024a); Hasan et al. (2024); Wang et al. (2022b); Marshall et al. (2015). Traditionally, machine learning models including support vector machines Marshall et al. (2014, 2017); Pereira et al. (2020), convolutional neural networks Marshall et al. (2017); Zhang et al. (2016) and logistic regression Millard et al. (2016); Marshall et al. (2020) are used for label prediction. More recently, some work employs transformers Dias et al. (2025) and LLMs to determine risk-of-bias levels Lai et al. (2024b); Pitre et al. (2023); Šuster et al. (2024). While RoBBR’s main task is similar to existing work, its novel contribution is the introduction of two new subtasks: support sentence retrieval (SSR) and support judgment selection (SJS). These subtasks identify crucial intermediate steps in risk-of-bias labeling and respectively measure the information retrieval and reasoning capabilities of LLMs when they assess risk-of-bias.
Support Sentences in Risk-of-Bias Labeling
Previous studies Marshall et al. (2016); Dias et al. (2025) attempt to extract supporting sentences from reviewers’ support judgments using direct quotes Marshall et al. (2016) or semantic embeddings Dias et al. (2025). However, these techniques falter because support judgments are free-form and frequently omit direct quotes, while semantic embeddings are imprecise and incomprehensive (See Table 2). This makes sentences extracted by such methods unsuitable for a rigorous benchmark like our Support Sentence Retrieval (SSR) subtask, which tests an LLM’s capability in retrieving support sentences. To address this, in Section 4.3, our annotation pipeline aligns all aspects (such as paraphrases, synthesized and summarized information) from support judgment to sentences in biomedical studies, which would more reliably and comprehensively extract support sentences that are used to form the novel SSR subtask.
LLM Extraction from Systematic Reviews
In recent years, LLMs have been used to extract data from systematic reviews, such as experiment evidence Sun et al. (2024); Gartlehner et al. (2023), summaries Wang et al. (2022a); Wallace et al. (2020), quality of evidence Suster et al. (2023) and PICO Tang et al. (2024). In our work, aligning the human support judgment to various sentences in biomedical studies is a more complex and difficult "data extraction" task, as noted by Dias et al., which necessitates the aspect decomposition and aspect & sentence annotation components in our annotation pipeline.
@@ 4 Benchmark Development
4.1 Benchmark Statistics and Data Sources
Cochrane Cochrane Non-Cochrane
Task Train Test Test
Retrieval (SSR) n=235 n=313 N/A
Selection (SJS) n=346 n=465 N/A
Risk-of-Bias (Main) n=774 n=906 n=2,489
Table 1: RoBBR Benchmark Statistics.
There are two prevailing risk-of-bias assessment guidelines, RoB1 Higgins and Green (2011) and the updated RoB2 Sterne et al. (2019). As of 2025, many recently published and updated systematic reviews still use RoB1, for example, Miao et al. (2025); Mehrholz et al. (2025). Since the RoB2 guidelines do not require support judgments that are necessary for our subtasks, our subtasks only include systematic reviews that follow RoB1. For the main task, we include systematic reviews that follow either set of guidelines.
Since there are hundreds of bias names and definitions, such as "incomplete adverse event reporting (reporting bias)" and "selective outcome reporting (reporting bias)", for clarity of presentation, we manually group them into 6 primary categories: "selection", "attrition", "performance", "detection", "reporting" and "deviation" (unique to RoB2) using two approaches. In some cases, bias names already include broader category labels (e.g., "random sequence generation (selection bias)"). When this happens, we follow these broader category labels. When bias names lack such broader category labels, we manually identify an equivalent bias name that has a category label and use this label. For example, the bias name "bias arising from the randomization process" is equivalent to "random sequence generation (selection bias)," so we use "selection bias" as its category label. A bias name belongs to multiple categories when these categories are explicitly included in the bias name. For example, a bias called "blinding for adverse event (performance and detection bias)" is categorized into both the "performance bias" category and the "detection bias" category. See Table 12 for examples of how bias names are categorized.
During evaluation, models would see the actual bias name and definition used in systematic reviews. Bias definitions are sourced from systematic reviews themselves and official guidelines like Higgins and Green (2011); Practice and of Care (EPOC); Sterne et al. (2014).
For data diversity, we include both systematic reviews that are published in Cochrane (denoted as Cochrane Train/Test) and that are published elsewhere (denoted as Non-Cochrane Test [^1] Empirically, Llama3 fine-tuned on Cochrane train set also perform well on Non-Cochrane test set, so we did not create a separate Non-Cochrane train set.). Non-Cochrane reviews do not have support judgments and cannot form our subtasks.
RoBBR includes 58 Cochrane reviews that assess 204 papers, and 496 Non-Cochrane reviews that assess 496 papers. Appendix A.3 show details on token length, bias and label distributions. See Table 1 for statistics of the three tasks in our benchmark.
4.2 Main Task: Risk-of-Bias Determination
The main task evaluates a model’s ability to label the risk-of-bias of a biomedical study as high, low, or unclear/some concern. The input to the model is exactly everything that a human expert would see, including the biomedical paper, study characteristics (i.e. PICO), the objective/topic of the systematic review, one specific bias name and definition, and risk level definition. See the top-left panel of Figure 1 for a visualization of input. The same input would be used for the two subtasks as described in section 4.4, 4.5. To visualize the task, see Figure 3.
Figure 3: Main Task: Risk-of-Bias Determination. The goal of the main task is to provide an assessment of the paper’s risk of specific biases.
The decision-making process of risk-of-bias determination is highly complex and involves multiple intermediate steps: retrieve support sentences from the biomedical paper, synthesize evidence from support sentences into a support judgment, and follow the guidelines and definitions of bias and risk level to give a final determination.
4.3 Automatic Annotation Pipeline
Support judgments contain a wealth of information, including evidence, data, reasoning, and commentaries written by expert reviewers. The goal of support judgment is to explain why reviewers would give this risk-of-bias rating. Therefore, support judgments mirror and shed light on the decision-making process of the reviewer.
It is difficult to find all support sentences in the study that form the basis of a support judgment. As shown in Figure 2, support judgments can have many aspects. While some aspects are direct quotes from the study, many other aspects are paraphrases from the study, or deeper analyses that synthesize information from multiple parts of the study.
To reverse engineer a support judgment, we find that no lexicon-based or embedding models can find all support sentences accurately (see Section 4.3.1 and Table 2). To address this, our proposed annotation pipeline has two main components: Aspect Decomposition and Aspect & Sentence Annotation. See Figure 2 for a visualization. GPT-4 (gpt-4-0125) is used by our annotation pipeline. All prompts used in our annotation pipeline are in Appendix D.
Aspect Decomposition:
To reduce the complexity of the reverse engineering process, we decompose the support judgment into distinct, non-overlapping pieces of information, each focusing on a specific aspect of the original support judgment. This is accomplished by careful prompt engineering with GPT-4. We manually validated that all decomposed aspects from all support judgments are of high quality.
Aspect & Sentence Annotation:
To further reduce complexity and instability introduced by long context input to LLMs, we only ask GPT-4 to determine if one sentence from the study covers one aspect.
Definition 1.
A sentence $s_{j}$ in a paper covers an aspect $f_{i}$ if it satisfies these two criteria:
1: The content of the sentence conveys the majority of the aspect’s information.
2: Any information about the aspect not directly stated in the sentence can be reasonably inferred from the surrounding context.
We aggregate these annotations into the Aspect-to-Sentence Mapping: For each (study, support judgment) triplet, a sentence is mapped to several aspects (e.g., 0, 1, 2, or more). Most sentences are not mapped to any aspect because they do not cover them. A few sentences can even be mapped to more than one aspect because aspect-to-sentence mapping is a many-to-many relationship.
The total cost of our annotation pipeline via OpenAI API is around \$1,000 with aspect decomposition and initial filtering costing less than \$100. Next, we evaluate the quality of GPT-4 annotation.
Metrics Human & Human Human & GPT p-value